Much progress has been made towards the BSM Phase II network, but much is yet to come!
BSM_Berry. This system is off-line until we have a replacement camera.
BSM_Hamren. We've fixed the roof controller by replacing its communication cable, and have added a ZWO electronic focuser. The system is fully operational, though we intend to replace the computer soon.
BSM_NH2. Fully operational. The new E180 astrograph has arrived, and Gary Walker is making a mounting plate. Viris-permitting, we'll start the upgrade process in a few weeks.
BSM_NM. Operational with the new E180 and ZWO ASI183mm-pro CMOS camera. We're still tweaking a couple of parameters to improve efficiency.
BSM_S. The declination drive failed. Peter has a new motor, so hopefully the system will be operational soon.
BSM_TX. Close to operation, but we're still having filter wheel problems that we hope to solve shortly.
OC61. Fully operational.
SRO. Fully operational.
TMO61. Engineering use only. We are in the process of replacing the camera with a QHY600, but are waiting for backordered filters.